[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Thu Oct 05, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Thu Oct 5 19:29:19 UTC 2017


HSU Shift Log for Thu Oct 05 2017 15:29:12 GMT-0400 (EDT)

Shift: H

Forecaster: E Blake

Technical Issues Summary:

ATCF issues:

N-AWIPS issues: Doug trying to fix mysterious GTWO launching issue on the  
Pacific side. It worked at 18z so maybe he succeeded. Also the ASCAT  
vectors are still showing up as circles on Tilefish.

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues: GOES-16 request needs to be re-entered after 03Z if  
the hurricane watch is issued for US, - only runs until 10Z because of  
outage.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20171005/c8e0f88b/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list